# Stochastic Multi-round Submodular Optimization with Budget

### v1

@@ 2.5.2 Stochastic Probing
The multi-round adaptive variant of this problem, that we call multi-round binary submodular stochastic probing (MR-BSS) assumes that the probabilities and the submodular set functions are time specific (i.e., we have a probability distribution $p_{t}$ and a function $g_{t}$ for any round $t$), the budget (i.e., the number of items we can select) can be adaptively allocated over all possible rounds, and the goal is to maximize the sum of the values achieved by all functions $g_{t}$. In the adaptive variant of the problem, one considers adaptive multi-round policies that adaptively select the items at each round. Furthermore, we assume that there are two quantities $\lambda,\Lambda>0$ that are polynomial in the input instance such that: (i) the increment caused to $g_{t}$, when probing a new item $v$, is upper bounded by $\Lambda$; (ii) there exists one round $t$ and an item $v$ such that the expected value of $g_{t}$ when probing item $v$ at round $t$ is at least $\lambda$.
@@
All desired properties of $f_{t}$ follows if $g_{t}$ also enjoys the same properties by definition.
@@ 3 Optimal Algorithm
In this section we first describe a single-round optimization problem of interest, and how its solution is used to provide a dynamic programming algorithm that will return the optimal policy in polynomial time as long as the single-round optimization problem can be efficiently addressed. Then, we will discuss a few of cases in which the single-round optimization problem can be efficiently addressed. Finally, we will highlight the limits of this dynamic programming framework, by discussing when it can be extended to provide a good approximation of the optimal policy, even if the single-round optimization problem can only be approximated.
@@ 3.1 A Dynamic Programming Algorithm based on Single-round Optimization
Let us fix a round $t$ and a budget $b$ available. Given $j\in\llbracket 0,b\rrbracket$, let $R(t+1,j)$ denote the optimal expected value of $\sum_{t^{\prime}=t+1}^{T}f_{t^{\prime}}$ achievable in rounds $t+1,\ldots,T$ by a multi-round policy $\pi_{\geq t+1}$ restricted to such rounds only; we observe that $R(t+1,j)=0$ for $t=T$. We assume that $R(t+1,k)$ has been already determined for any $k\in\llbracket 0,b\rrbracket$. The single-round optimization problem at round $t$ asks to find a policy $\pi_{t}$ at round $t$ that adaptively computes the budget $\bm{i}\in\llbracket 0,b\rrbracket$ to be allocated at round $t$ in such a way that the expected value
$\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-\bm{i})]$ (1)
of state function $f_{t}$ plus $R(t+1,b-\bm{i})$ is maximized, where the random budget $\bm{i}$ used at round $t$ depends on the chosen policy $\pi_{t}$ and the state $\bm{\eta}_{t}$.
@@
$R(t,b)=\begin{cases}0,&\text{ if }t=T+1\\ \max_{\pi_{t}}\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-\bm{i})],&\text{ if }t\in\llbracket 0,T\rrbracket.\end{cases}$ (2)
Suppose that there is an algorithm for the single-round optimization problem described above, that is, by resorting to the recursive relation (2), we can efficiently compute $R(t,b)$ as a function of $R(t+1,j)$ for any $j\in\llbracket 0,b\rrbracket$, together with the optimal single-round policy. Such algorithm can be used as a sub-routine for a more general dynamic programming algorithm that computes an optimal policy for the general SBMSm problem.
Specifically, the dynamic programming algorithm first sets $R(T+1,b)=0$ for every $b\in\llbracket 0,B\rrbracket$. By resorting to the recursive relation (2), we can compute $R(T,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(T+1,j)$ with $j\in\llbracket 0,b\rrbracket$. Iteratively, going from $t=T-1$ to $t=1$, we can compute $R(t,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(t+1,j)$ with $j\in\llbracket 0,b\rrbracket$. Finally, we observe that $R(1,B)$ is the optimal value for the SBMSm problem. Thus, starting from round $t=1$ and budget $B$, we can reconstruct the optimal multi-round policy achieving the optimal value (this can be done by resorting to standard dynamic programming techniques).
@@ 3.2 How to Solve Single-round Optimization Efficiently?
In the above dynamic-programming algorithm, we assumed the existence of a sub-routine that solves the single-round optimization problem, i.e., that allows to compute recurrence (2). Unfortunately, the considered single-round problem can be easily reduced from the submodular monotone optimization problem, that is NP-hard [19]. To deal with the intractability of the single-round optimization problem, we will describe an algorithm that is in general inefficient, but that can be run in polynomial time in a few cases.
@@
To compute a single-round policy $\pi_{t}$ that computes $R(t,b)$ for any round $t$ and budget $b$, it is useful to describe the single-round optimization problem at round $t$ in terms of a suitable two-player extensive form game between two players [27]: (i) the policy $\pi_{t}$, that decides, at each step, either to interrupt the game and collecting a reward or to continue by selecting another item; (ii) the nature $\mathcal{N}$, that reveals a partial state after each item selection.
@@
Case $d\in\{0,1\}$: The root $u^{*}=u_{0}$ of $\mathcal{T}_{t}$, that belongs to $\mathcal{T}_{t}^{o}$ (as its depth is 00), is the initial configuration of the game. The root has $n+1$ outgoing edges, where each edge $(u_{0},u_{1})$ has a label $a=L(u_{0},u_{1})$ corresponding to the action $a$ that $\pi_{t}$ can perform at step $1$. In particular, either the action/label $a$ is a string STOP that will denote the termination of the single-round policy, or $a$ is an item $v_{1}\in V$ that $\pi_{t}$ can select at step $1$. Given the action $a$ chosen by $\pi_{t}$ at step $1$, the new configuration of the game is obtained by following the corresponding labeled edge (i.e., the edge $(u_{0},u_{1})$ such that $a=L(u_{0},u_{1})$), and it corresponds to a node $u_{1}$ at depth $1$. If $a=$STOP then the reached node $u_{1}$ is a leaf and the game ends. In such a case, the reward of $\pi_{t}$ is $R(t+1,b)$, as $\pi_{t}$ did not select any node and the whole budget $b$ can be used in rounds $t+1,\ldots,T$. Otherwise, the reached node $u_{1}$ has as many outgoing edges as partial states $\langle\{v_{1}\}\rangle$ that can be observed after selecting the item $v_{1}$, and each outgoing edge $(u_{1},u_{2})$ is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1}\}\rangle$, drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1}\}\rangle$ (i.e., observable after selecting $v_{1}$). Then, the new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state (i.e., the edge $(u_{1},u_{2})$ such that $\langle\{v_{1}\}\rangle=L(u_{1},u_{2})$), and it corresponds to a node $u_{2}$ at depth $2$, that belongs to $\mathcal{T}_{t}^{o}$.
Case $d\in\{2,3\}$: Now, $u_{2}$ has $n$ outgoing edges, again labeled with the actions $a$ that $\pi_{t}$ can perform at step $2$. In particular, either $a$=STOP, or $a$ is equal to an item $v_{2}\in V\setminus\{v_{1}\}$ that $\pi_{t}$ can select at step $2$. Given the action $a$ chosen by $\pi_{t}$ at step $2$, the new configuration of the game is obtained by following the corresponding labeled edge (i.e., the edge $(u_{2},u_{3})$ such that $a=L(u_{2},u_{3})$), and it corresponds to a node $u_{3}$ at depth $3$. If $a=$STOP then $u_{3}$ is a leaf, the game ends and the reward for $\pi_{t}$ is the expected value of $f_{t}$ constrained by the observation of partial state $\langle\{v_{1}\}\rangle$, plus the optimal expected value $R(t+1,b-1)$ achievable in the subsequent rounds $t+1,\ldots,T$ by using the residual budget (i.e., $b-1$ units). Otherwise, the reached node $u_{3}$ has as many outgoing edges $(u_{3},u_{4})$ as partial states $\langle\{v_{1},v_{2}\}\rangle$ that can be observed after selecting $v_{2}$ (constrained by the previous observation of $\langle\{v_{1}\}\rangle$), and each outgoing edge is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1},v_{2}\}\rangle$, drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1},v_{2}\}\rangle$, conditioned by the observation of $\langle v_{1}\rangle$. The new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state, thus reaching a node $u_{4}$ at depth $4$, that belongs to $\mathcal{T}_{t}^{o}$.
@@
Case $d\in\{2(i-1),2i-1\}$: Iteratively, for any $i\in[n+1]$, each node $u_{d}\in\mathcal{T}_{t}$ at depth $d=2(i-1)$ corresponds to a possible outcome of policy $\pi_{t}$ obtained at the end of step $i-1$, where $i-1$ items $v_{1},\ldots,v_{i-1}$ have been already selected and a partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$ has been observed. $u_{d}$ has $n-i+2$ outgoing edges $(u_{d},u_{d+1})$, each one labeled with the action $a=L(u_{d},u_{d+1})$ that $\pi_{t}$ can performs at step $i$: either $a$=STOP, or $a$ is equal to an item $v_{i}\in V\setminus\{v_{1},\ldots,v_{i-1}\}$ that $\pi_{t}$ can select at step $i$; we observe that, for $i=n+1$, there is a unique outgoing edge labeled with action $a=$STOP. Again, given the action $a$ chosen by $\pi_{t}$ at step $i$, the new configuration of the game is obtained by following the corresponding labeled edge and reaching a node $u_{d+1}$ at depth $d+1$. If $a=$STOP, then $u_{d+1}$ is a leaf, the game ends and the profit for $\pi_{t}$ is the expected value of $f_{t}$ constrained by the observation of partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$, plus the optimal expected value $R(t+1,b-i+1)$ achievable in the subsequent rounds by using the residual budget (i.e., $b-i+1$ units). Otherwise, $u_{d+1}$ has as many outgoing edges as partial states $\langle\{v_{1},\ldots,v_{i}\}\rangle$ that can be observed after selecting $v_{i}$ (constrained by the previous observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$), and each outgoing edge is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1},\ldots,v_{i}\}\rangle$ drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1},\ldots,v_{i}\}\rangle$, conditioned by the observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$. The new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state, and it corresponds to a node $u_{d+1}$ at depth $d+1$.
@@
Given a node $u\in\mathcal{T}_{t}^{o}$, let $\langle S(u)\rangle$ denote the partial state observed when the configuration of the game is determined by $u$. We have that the above surjective map between nodes of $\mathcal{T}_{t}^{o}$ and partial states defines an equivalence between the tree-based game and the problem of finding a policy $\pi_{t}$ that maximizes (1). Indeed, let $a^{*}(u)$ denote the action that maximizes the expected value of the reward collected by $\pi_{t}$, assuming that the actual configuration of the game is $u$, and let $R(u)$ denote the resulting optimal expected value. We have that $a^{*}(u)$ is also the action that $\pi_{t}$ must follow for any partial state $\langle S(u)\rangle$ in order to maximize (1). Then, denoting the root of $\mathcal{T}_{t}$ by $u^{*}$, we have that $R(u^{*})$ coincides with $R(t,b)$, that is the optimal value of the single-round optimization problem at round $t$ with budget $b$.
We observe that, by following standard approaches to find the optimal strategy of a player in extensive-form games [27], $a^{*}(u)$ and $R(u)$ can be computed for any $u\in\mathcal{T}_{t}^{o}$ by recursively exploring the tree $\mathcal{T}$. By the equivalence bewteen the tree-based game and the single-round optimization problem, the optimal strategy for the former can be easily turned into an optimal solution for the latter.
@@
By using the above tree-based algorithmic procedure, the single-round optimization problem can be solved efficiently whenever the game tree ${\cal T}_{t}$ contains a number of nodes that is polynomial in the size of the input instance of the general SBMSm problem. To this aim, in the proposition below, we relate the size of ${\cal T}_{t}$ to the number of items and states, and we next show that there may be practical settings in which this is not rare.
@@
By exploiting the structure of $\mathcal{T}_{t}$, we have that each non-leaf node $u\in\mathcal{T}_{t}^{o}$ has exactly one child node that is a leaf. Thus, the leaves of $\mathcal{T}$ are at least as numerous as the nodes of $\mathcal{T}_{t}^{o}$, that is, $|\mathcal{T}_{t}^{o}|\leq z$. As the father of each node of ${\cal T}^{e}_{t}$ belongs to ${\cal T}^{o}_{t}$, we necessarily have $|{\cal T}^{e}_{t}|\geq|{\cal T}^{o}_{t}|$. By putting the above inequalities together, we get $|{\cal T}_{t}|=|{\cal T}^{o}_{t}|+|{\cal T}^{e}_{t}|\leq 2|{\cal T}^{e}_{t}|\leq 2z$, and this shows (3).
@@
Let $G$ be the function that maps each sequence $s$ of distinct items in $V$ and partial state state $\eta_{t}\in H$, to the leaf $u=G(s,\eta_{t})$ of $\mathcal{T}_{t}$ associated with the outcome of $\pi_{t}$ obtained when the realized state is $\eta_{t}$ and $s$ is the sequence of items selected by $\pi_{t}$ at round $t$. We observe that, by the structure of $\mathcal{T}_{t}$, function $G$ is surjective. Thus, as the number of sequences of distinct items in $V$ is at most $2^{n}n!$ (at most $n!$ orderings for each of the $2^{n}$ subsets of $V$), the number of leaves is at most $2^{n}n!|H|$, and this shows (4). By putting (3) and (4) together, the claim of the proposition follows. ∎
If $n!$ and $|H|$ are polynomial in the input instance of SBMSm, then $2^{n}n!|H|$ is polynomial, too. In such a case, by Proposition 1, the number of nodes of ${\cal T}_{t}$ is polynomial. Thus, the tree-based algorithm for the single-round optimization problem can be run in polynomial time, and consequently so can then the dynamic programming algorithm for SBMSm.
Note that $n!$ is polynomial if, for instance, the number $n$ of items is very small, e.g., constant. Such a limitation is not so rare in several scenarios of interest, as it often occurs that in large populations only few items are really relevant (e.g., among the many friends on social networks only few of them are able to affect our opinions, or among the many potential job candidates only few effectively satisfy all the requirements at which one may be interested in).
The limitation on the number of states $|H|$ occurs in many settings where the number of observable events is not so high. For instance, it could be the case of certain influence maximization problems where edge probabilities are correlated, and depends on a few possible stochastic events. Furthermore, it often occurs that, if the number of items is constant, the number of states $|H|$ is constant, too. For instance, in both influence maximization and stochastic probing, we have that the number of states, even if it grows exponentially in the number of items, it depends on such quantity only. Therefore, if the number of items is constant, the number of states will be constant, too.
@@
While it would be possible to fix our dynamic programming framework in order to do not cause such an explosion in the approximation guarantee, and to provide an $\left(\frac{1}{2}-\varepsilon\right)$-approximation to the optimal multi-round policy, we next prove that the same bound can be actually achieved by a simpler and less computationally expensive greedy algorithm.
4 Semi-Adaptive Approximation Algorithms
We now consider a semi-adaptive policy based on the greedy approach, that works as follows: it first determines how to distribute the budget $B$ over all rounds by computing a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$, so that $b_{t}$ is the budget assigned to round $t$; then, for each round $t\in[T]$, it runs an adaptive policy that greedily selects $b_{t}$ items to be observed in round $t$.
It is worth to notice that an optimal solution for SBMSm, in general, dynamically allocates the budget after each item selection. Instead our semi-adaptive approximate policy, at the expense of a small loss in the optimality guarantee, allocates all the budget to the rounds at the beginning.

### v2

@@ 2.5.2 Stochastic Probing
The multi-round adaptive variant of this problem, denoted as multi-round binary submodular stochastic probing (MR-BSS), assumes that the probabilities and the submodular set functions are time/round specific (i.e., we have a probability distribution $p_{t}$ and a function $g_{t}$ for any round $t$), the budget (i.e., the number of items we can select) can be adaptively allocated over all possible rounds, and the goal is to maximize the sum of the values achieved by all functions $g_{t}$. In the adaptive variant of the problem, one considers adaptive multi-round policies that adaptively select the items at each round. Furthermore, we assume that there are two quantities $\lambda,\Lambda>0$ that are polynomial in the input instance such that: (i) the increment caused to $g_{t}$, when probing a new item $v$, is upper bounded by $\Lambda$; (ii) there exists one round $t$ and an item $v$ such that the expected value of $g_{t}$ when probing item $v$ at round $t$ is at least $\lambda$.
@@
The polynomial boundedness of $f_{t}$ holds by using the same values of $\lambda$ and $\Lambda$ defined for $g_{t}$. The monotonicity and the adaptive submodularity of $f_{t}$ immediately follow from the monotonicity and the submodularity of $g_{t}$, respectively.
@@ 3 Optimal Algorithm
In this section we first describe a single-round optimization problem of interest, and how its solution is used to provide a dynamic programming algorithm that will return the optimal policy in polynomial time as long as the single-round optimization problem can be efficiently addressed. Then, we will discuss a few of cases in which the single-round optimization problem can be efficiently addressed.
These results will imply that, if the number of states $|H|$ is polynomially bounded [^3] In the model definition, we already observed that $|H|$ might be exponential (e.g., for influence maximization problems). in the size $|I|$ of the input instance, SBMSm is fixed-parameter tractable (FTP) [22] w.r.t. the number of items $n$, that is, the complexity of the (optimal) dynamic programming algorithm is $O(F(n)\cdot|I|^{p})$ for some computable function $F$ and $p>0$.
Finally, we will highlight the limits of this dynamic programming framework, by discussing when it can be extended to provide a good approximation of the optimal policy, even if the single-round optimization problem can only be approximated.
@@ 3.1 A Dynamic Programming Algorithm based on Single-round Optimization
Let us fix a round $t$ and a budget $b$ available. Given $i\in\llbracket 0,\min\{n,b\}\rrbracket$, let $R(t+1,i)$ denote the optimal expected value of $\sum_{t^{\prime}=t+1}^{T}f_{t^{\prime}}$ achievable in rounds $t+1,\ldots,T$ by a multi-round policy $\pi_{\geq t+1}$ that is restricted to such rounds only and uses at most $i$ units of budget; we observe that $R(t+1,i)=0$ for $t=T$. We assume that $R(t+1,i)$ has been already determined for any $i\in\llbracket 0,b\rrbracket$. The single-round maximization problem at round $t$ and budget $b$ ($\text{SRm}(t,b)$) asks to find a policy $\pi_{t}$ at round $t$ that adaptively computes the budget $i(\pi_{t},\bm{\eta}_{t})\in\llbracket 0,b\rrbracket$ to be allocated at round $t$ in such a way that the expected value
$\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-i(\pi_{t},\bm{\eta}_{t}))]$ (1)
of objective function $f_{t}$ plus $R(t+1,b-i(\pi_{t},\bm{\eta}_{t}))$ is maximized, where the random budget $i(\pi_{t},\bm{\eta}_{t})$ used at round $t$ depends on the chosen policy $\pi_{t}$ and the state $\bm{\eta}_{t}$.
@@
$R(t,b)=\begin{cases}0,&\text{ if }t=T+1\\ \max_{\pi_{t}}\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-i(\pi_{t},\bm{\eta}_{t}))],&\text{ if }t\in\llbracket 0,T\rrbracket.\end{cases}$ (2)
Suppose that there is an algorithm for $\text{SRm}(t,b)$, that is, by resorting to the recursive relation (2), we can efficiently compute $R(t,b)$ as a function of $R(t+1,i)$ for any $i\in\llbracket 0,b\rrbracket$, together with the optimal single-round policy. Such algorithm can be used as a sub-routine for a more general dynamic programming algorithm that computes an optimal policy for the general SBMSm problem.
Specifically, the dynamic programming algorithm first sets $R(T+1,b)=0$ for every $b\in\llbracket 0,B\rrbracket$. By resorting to the recursive relation (2), we can compute $R(T,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(T+1,b-i)$ with $i\in\llbracket 0,\min\{n,b\}\rrbracket$. Iteratively, going from $t=T-1$ to $t=1$, we can compute $R(t,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(t+1,b-i)$ with $i\in\llbracket 0,\min\{n,b\}\rrbracket$. Finally, we observe that $R(1,B)$ is the optimal value for the SBMSm problem. Thus, starting from round $t=1$ and budget $B$, we can reconstruct the optimal multi-round policy achieving the optimal value (this can be done by resorting to standard dynamic programming techniques).
@@ 3.2 How to Solve Single-round Optimization Efficiently?
In the above dynamic-programming algorithm, we assumed the existence of a sub-routine that solves $\text{SRm}(t,b)$ for any round $t$ and budget $b$, i.e., that allows to compute recurrence (2). Unfortunately, the $\text{SRm}(t,b)$ problem can be easily reduced from the submodular monotone optimization problem, that is NP-hard [23]. To deal with the intractability of the single-round optimization problem, we will describe an algorithm that is in general inefficient, but that can be run in polynomial time in a few cases.
@@
To design a single-round policy $\pi_{t}$ that computes $R(t,b)$ for any round $t$ and budget $b$, it is useful to describe the $\text{SRm}(t,b)$ problem at round $t$ in terms of a suitable two-player extensive form game between two players [33]: (i) the policy $\pi_{t}$, that decides, at each step, either to interrupt the game and collecting a reward or to continue by selecting another item; (ii) the nature $\mathcal{N}$, that reveals a partial state after each item selection.
@@
Base Case ($d\in\{0,1\}$): The root $u^{*}=u_{0}$ of $\mathcal{T}_{t}$, that belongs to $\mathcal{T}_{t}^{o}$ (as its depth is 00), is the initial configuration of the game. The root has $n+1$ outgoing edges, where each edge $(u_{0},u_{1})$ has a label $a=L(u_{0},u_{1})$ corresponding to the action $a$ that $\pi_{t}$ can perform at step $1$. In particular, either the action/label $a$ is a string STOP that will denote the termination of the single-round policy, or $a$ is an item $v_{1}\in V$ that $\pi_{t}$ can select at step $1$. Given the action $a$ chosen by $\pi_{t}$ at step $1$, the new configuration of the game is obtained by following the corresponding labeled edge (i.e., the edge $(u_{0},u_{1})$ such that $a=L(u_{0},u_{1})$), and it corresponds to a node $u_{1}$ at depth $1$. If $a=$STOP then the reached node $u_{1}$ is a leaf and the game ends. In such a case, the reward of $\pi_{t}$ is $R(t+1,b)$, as $\pi_{t}$ did not select any node and the whole budget $b$ can be used in rounds $t+1,\ldots,T$. Otherwise, the reached node $u_{1}$ has as many outgoing edges as partial states $\langle\{v_{1}\}\rangle$ that can be observed after selecting the item $v_{1}$, and each outgoing edge $(u_{1},u_{2})$ is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1}\}\rangle$, drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1}\}\rangle$ (i.e., observable after selecting $v_{1}$). Then, the new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state (i.e., the edge $(u_{1},u_{2})$ such that $\langle\{v_{1}\}\rangle=L(u_{1},u_{2})$), and it corresponds to a node $u_{2}$ at depth $2$, that belongs to $\mathcal{T}_{t}^{o}$.
@@
General Case ($d\in\{2(i-1),2i-1\}$): Iteratively, for any $i\in[n+1]$, each node $u_{d}\in\mathcal{T}_{t}$ at depth $d=2(i-1)$ corresponds to a possible outcome of policy $\pi_{t}$ obtained at the end of step $i-1$, where $i-1$ items $v_{1},\ldots,v_{i-1}$ have been already selected and a partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$ has been observed. $u_{d}$ has $n-i+2$ outgoing edges $(u_{d},u_{d+1})$, each one labeled with the action $a=L(u_{d},u_{d+1})$ that $\pi_{t}$ can performs at step $i$: either $a$=STOP, or $a$ is equal to an item $v_{i}\in V\setminus\{v_{1},\ldots,v_{i-1}\}$ that $\pi_{t}$ can select at step $i$; we observe that, for $i=n+1$, there is a unique outgoing edge labeled with action $a=$STOP. Again, given the action $a$ chosen by $\pi_{t}$ at step $i$, the new configuration of the game is obtained by following the corresponding labeled edge and reaching a node $u_{d+1}$ at depth $d+1$. If $a=$STOP, then $u_{d+1}$ is a leaf, the game ends and the profit for $\pi_{t}$ is the expected value of $f_{t}$ constrained by the observation of partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$, plus the optimal expected value $R(t+1,b-i+1)$ achievable in the subsequent rounds by using the residual budget (i.e., $b-i+1$ units). Otherwise, $u_{d+1}$ has as many outgoing edges as partial states $\langle\{v_{1},\ldots,v_{i}\}\rangle$ that can be observed after selecting $v_{i}$ (constrained by the previous observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$), and each outgoing edge is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1},\ldots,v_{i}\}\rangle$ drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1},\ldots,v_{i}\}\rangle$, conditioned by the observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$. The new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state, and it corresponds to a node $u_{d+1}$ at depth $d+1$.
@@
Given a node $u\in\mathcal{T}_{t}^{o}$, let $\langle S(u)\rangle$ denote the partial state observed when the configuration of the game is determined by $u$. We have that the above surjective map between nodes of $\mathcal{T}_{t}^{o}$ and partial states defines an equivalence between the tree-based game and the problem of finding a policy $\pi_{t}$ that maximizes (1). Indeed, let $a^{*}(u)$ denote the action that maximizes the expected value of the reward collected by $\pi_{t}$, assuming that the actual configuration of the game is $u$, and let $R(u)$ denote the resulting optimal expected value. We have that $a^{*}(u)$ is also the action that $\pi_{t}$ must follow for any partial state $\langle S(u)\rangle$ in order to maximize (1). Then, denoting the root of $\mathcal{T}_{t}$ by $u^{*}$, we have that $R(u^{*})$ coincides with $R(t,b)$, that is the optimal value of the $\text{SRm}(t,b)$ problem.
We observe that, by following standard approaches to find the optimal strategy of a player in extensive-form games [33], $a^{*}(u)$ and $R(u)$ can be computed for any $u\in\mathcal{T}_{t}^{o}$ by recursively exploring the tree $\mathcal{T}$. By the equivalence between the tree-based game and the single-round maximization problem, the optimal strategy for the former can be easily turned into an optimal solution for the latter.
@@
The tree-based algorithmic procedure to compute an optimal policy for $\text{SRm}(t,b)$ can be run in linear time w.r.t. the size of the game tree $\mathcal{T}_{t}$, as it simply requires to recursively visit all the nodes in tree. Thus, $\text{SRm}(t,b)$ can be solved efficiently whenever the game tree ${\cal T}_{t}$ contains a number of nodes that is polynomial in the size of the input instance. To this aim, in the proposition below, we relate the size of ${\cal T}_{t}$ to the number of items and states, and we next show that there may be practical settings in which this is not rare.
@@
By exploiting the structure of $\mathcal{T}_{t}$, we have that each non-leaf node $u\in\mathcal{T}_{t}$ has exactly one child node that is a leaf. Then, denoting as $y$ the number of non-leaf nodes of $\mathcal{T}_{t}$, we have $y\leq z$. Thus, we have $|{\cal T}_{t}|=y+z\leq 2z$, and then (3) follows.
@@
Let $G$ be the function that maps each sequence $s$ of distinct items in $V$ and partial state state $\eta_{t}\in H$, to the leaf $u=G(s,\eta_{t})$ of $\mathcal{T}_{t}$ associated with the outcome of $\pi_{t}$ obtained when the realized state is $\eta_{t}$ and $s$ is the sequence of items selected by $\pi_{t}$ at round $t$. We observe that, by the structure of $\mathcal{T}_{t}$, function $G$ is surjective. Thus, as the number of sequences of distinct items in $V$ is at most $2^{n}n!$ (at most $n!$ orderings for each of the $2^{n}$ subsets of $V$), we have that the number of leaves is at most $2^{n}n!|H|$, and this shows (4).
By putting (3) and (4) together, the claim of the proposition follows. ∎
Since the tree-based algorithm for $\text{SRm}(t,b)$ can be run in linear time w.r.t. to the size of $\mathcal{T}_{t}$ and, by Proposition 1, such size is $O(2^{n}n!|H|)$, we then have the following corollary:
Corollary 1.
$\text{SRm}(t,b)$ can be solved in $O(2^{n}n!|H|)$ time.
Furthermore, by observing that the dynamic programming algorithm for SBMSm executes $\text{SRm}(t,b)$ once for each pair round-budget $(t,b)\in[T]\times\llbracket 0,B\rrbracket$, i.e., $O(BT)\subseteq O(nT^{2})$ times, we have the following further corollary:
Corollary 2.
SBMSm can be solved in $O(2^{n}n!nT^{2}|H|)$ time.
If $n!$ and $|H|$ are polynomial in the input instance of SBMSm, then $2^{n}n!|H|$ is polynomial, too. Thus, by Corollaries 1 and 2, both $\text{SRm}(t,b)$ and SBMSm can be run in polynomial time in such case. Note that $n!$ is polynomial if, for instance, the number $n$ of items is very small, e.g., constant. Such a limitation is not so rare in several scenarios of interest, as it often occurs that in large populations only few items are really relevant (e.g., among the many friends on social networks only few of them are able to affect our opinions, or among the many potential job candidates only few effectively satisfy all the requirements at which one may be interested in). The limitation on the number of states $|H|$ occurs in many settings where the number of observable events is not so high. For instance, it could be the case of certain influence maximization problems where edge probabilities are correlated, and depends on a few possible stochastic events. Furthermore, it often occurs that, if the number of items is constant, the number of states $|H|$ is constant, too. For instance, in both influence maximization and stochastic probing, we have that the number of states, even if it grows exponentially in the number of items, it depends on such quantity only. Therefore, in such a case, if the number of items is constant, the number of states will be constant, too.
Finally, by exploiting again the complexity bounds provided in Corollaries 1 and 2, we have the following result on fixed-parameter-tractability:
Corollary 3.
Both $\text{SRm}(t,b)$ and SBMSb are fixed-parameter-tractable problems w.r.t. to $n$, if $|H|$ is polynomially bounded in the size of the input instance.
@@
While it would be possible to fix our dynamic programming framework in order to do not cause such an explosion in the approximation guarantee, and to provide an $\left(1-1/e-\epsilon\right)$-approximation to the optimal multi-round policy, we next prove that the same bound can be actually achieved by a simpler and less computationally expensive greedy algorithm.
4 Partially Adaptive Approximation Algorithm
In this section, we will design and analyze a partially adaptive policy as approximation algorithm for SBMSm. In general, a partially adaptive policy, in a first part, determines how to distribute the budget $B$ over all rounds by computing a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$, so that $b_{t}$ is the budget assigned to round $t$; this budget assignment is done in a non-adaptive way, that is, without knowing the particular realization of the stochastic environment, but only based on the knowledge of the input instance $I$. Subsequently, for each round $t\in[T]$, the policy adaptively selects $b_{t}$ items to be observed in round $t$; differently from the the first part, now the realization of the stochastic environment is partially revealed after each observation. It is worth to notice that an optimal solution for SBMSm, in general, dynamically allocates the budget after each item selection. Instead a partially adaptive policy allocates all the budget to the rounds at the beginning.
The specific partially adaptive policy we will consider is based on the greedy approach: (i) it first determines the budget allocation by repeatedly assigning each unit of budget to the rounds in a non-adaptive greedy way, and then (ii) greedily selects the items in an adaptive way (according to the budget previously allocated to each round). Despite partially adaptive policies are in general weaker than fully adaptive ones (because of the relaxation on budget adaptivity), we will show that our policy still guarantees a good approximation of the optimal adaptive value. In particular, we will show that its approximation factor is $1-1/e-\epsilon\approx 0.632-\epsilon$, where $e\approx 2.718$ denotes the Nepero number and $\epsilon>0$ is an arbitrarily small error due to some sampling procedures. Furthermore, as a consequence of the hardness result of [23] (holding for non-adaptive single-round submodular maximization), we have that SBMSm cannot approximated within a factor better than $1-1/e$ (unless P=NP). Thus, the approximation guaranteed by our partially adaptive policy cannot be substantially improved.